# Evaluating Code-Mixing in LLMs Across 18 Languages

### v1

@@ 5 Conclusion
This study introduces CodeMixEval, a comprehensive evaluation framework for assessing code-mixing performance in LLMs, spanning eight tasks and 18 languages. We also adopt GPT-4 Turbo for constructing synthetic code-mixed data to address data scarcity issues. Our findings show that while code-mixing challenges LLMs performance, improvements can be achieved through larger pre-training datasets, increased model scales, and few-shot learning. In the future, CodeMixEval holds great promise for evaluating the code-mixing capabilities of LLMs and inspiring further research in this area.
@@
While CodeMixEval provides a broad coverage of 18 languages, several limitations should be noted. First, both the synthetic and open-source datasets may contain quality inconsistencies despite rigorous filtering procedures. Second, our language coverage excludes some significant multilingual regions (e.g., Southeast Asia, Africa) due to resource constraints and the high computational costs associated with GPT-4 Turbo-based synthesis. These factors also limited the diversity of tasks we could include, potentially affecting the comprehensiveness of our evaluation.
Appendix A CodeMixEval vs. other benchmarks
@@ Appendix D Semantic Filtering using LaBSE
Table 5: LaBSE scores for synthetic code-mixed data across different language pairs. $L1$ indicates non-English languages. $L2$ denotes English. $CM$ denotes synthesized code-mixing data.
@@ Appendix E Model Aligned Filtering using GPT-4
Appendix F Prompts for Building CodeMixEval
@@ F.1 Synthesis Prompt
$<$word/phrase-->word/phrase$>$,
$<$word/phrase-->word/phrase$>$, …
@@ Appendix H Statistics of Synthetic Datasets
Lang. Size L1 / L2 / Other tokens Word-Level Semantic Model-Aligned
@@
Table 7: The statistics of synthesized datasets. The column $Lang.$ indicates the two languages code-mixed in the dataset. The column $Size$ indicates the size of the dataset. The column $L1/L2/Other$ $tokens$ shows token counts for the first language, the second language, and other language-independent tokens. In the column $Word$-$Level$, $M$ indicates the M-index, and $I$ indicates the I-index. In the column $Semantic$, $sim1$ represents the similarity between code-mixed text and the monolingual text in the first language, $sim2$ the similarity with the text in the second language, and $sim3$ the similarity between the monolingual texts in the first and second languages. In the $Model$-$Aligned$ column, $Co.$, $Na.$, and $Re.$ denote coherence, naturalness, and readability respectively.
Figure 6: Distribution of Additional synthetic code-mixing datasets in CodeMixEval.
Figure 7: Distribution of Additional synthetic code-mixing datasets in CodeMixEval.
@@ Appendix I Experiment Results of Collected Datasets
Named Entity Recognition (Accuracy)

### v2

@@ 5 Conclusion
This study introduces CodeMixBench, a comprehensive benchmark for evaluating code-mixing performance in LLMs, spanning eight tasks and 18 languages. We also adopt GPT-4 Turbo for constructing synthetic code-mixed data to address data scarcity issues. Our findings show that while code-mixing challenges LLMs performance, improvements can be achieved through larger pre-training datasets, increased model scales, and few-shot learning. In the future, CodeMixBench holds great promise for evaluating the code-mixing capabilities of LLMs and inspiring further research in this area.
@@
We introduce CodeMixBench, a collection of 22 synthetic datasets and 30 open-source datasets, each with potential quality issues. Our synthesis method generates large-scale code-mixed datasets with detailed filtering, but unexpected quality problems can still occur. Furthermore, our benchmark includes 18 languages, making it challenging to maintain consistent quality control. Furthermore, it could be promising to evaluate and mitigate the potential bias in code-mixing scenarios Ravfogel et al. (2020); Peng et al. (2025).
Acknowledgments
We would like to thank all anonymous reviewers for their insightful comments and feedback.
Appendix A CodeMixBench vs. other benchmarks
@@ Appendix D Semantic Filtering using LaBSE
Table 5: LaBSE scores for synthetic code-mixed data across different language pairs. $L1$ indicates non-English languages. $L2$ denotes English. CMCM denotes synthesized code-mixing data.
@@ Appendix E Model Aligned Filtering using GPT-4
Appendix F Prompts for Building CodeMixBench
@@ F.1 Synthesis Prompt
<<<word/phrase-->word/phrase>>>,
<<<word/phrase-->word/phrase>>>, …
@@ Appendix H Statistics of Synthetic Datasets
Lang. Size
L1 / L2 / Other
tokens
Word-Level Semantic Model-Aligned
@@
Table 7: The statistics of synthesized datasets. The column $Lang.$ indicates the two languages code-mixed in the dataset. The column SizeSize indicates the size of the dataset. The column $L1/L2/Other$ tokenstokens shows token counts for the first language, the second language, and other language-independent tokens. In the column WordWord-LevelLevel, $M$ indicates the M-index, and II indicates the I-index. In the column SemanticSemantic, $sim1$ represents the similarity between code-mixed text and the monolingual text in the first language, $sim2$ the similarity with the text in the second language, and $sim3$ the similarity between the monolingual texts in the first and second languages. In the ModelModel-AlignedAligned column, $Co.$, $Na.$, and $Re.$ denote coherence, naturalness, and readability respectively.
Figure 6: Distribution of Additional synthetic code-mixing datasets in CodeMixBench.
Figure 7: Distribution of Additional synthetic code-mixing datasets in CodeMixBench.
@@ Appendix I Experiment Results of Collected Datasets
Named Entity Recognition (F1)